INAF queue observing January-February 2014

Observers: E. Sani, F. Cusano, R. Carini, S. Bisogni
Telescope Operators: G. Bechetti, D. Gonzales-Huerta, S. Allanson
Support Astronomer: O. Khun, M. Edwards
Instrument: MODS, LBC

January 27-28

Night in shared time with AZ time. We start observations at 07:30 UT
At the beginning of observations we have thin cirrus above the orison (NO direction), low wind ans humidity at 26%.
N.B. The seeing from the GCS is overestimated by 10% because the MODs guider is out of focus.

Project_RUN Objects Filter ExpTime Notes
Standard Feige34 red grating - airmass 1.02. Default 5" slit + 1" slit
Gilli_4 J1048 Mask1 red grating 3.5h the last 1.5h have an average seeing out of specs

January 28-29

At the beginning of the twilight we have mostly clear sky (thin cirrus just above the NO horizon), with no wind and low humidity (15%).
During the first part of night the seeing is good, ranging between 0.8 and 1.0 on the DIMM. Also the humidity is very low (6%) with no wind
The blu-shutter remained closed during the OB execution. The SA fix the problem turning off and on the camera power from the engineering interface.
At 5.0UT we get problems with the blue channel collimation. We clean the active optics and commit manual corrections for coma.
At 9.0UT thin cirrus passing above N-O horizon.
The shutter problem happened more than once during the night.

Project_RUN Objects Filter ExpTime Notes
Battaglia_DDT IC10blob g-Sloan, i-Sloan 15min, 15 min COMPLETED
Battaglia_DDT IC10fore g-Sloan, i-Sloan 15min, 15 min COMPLETED
Battaglia_DDT IC10ext g-Sloan, i-Sloan 15min, 15 min COMPLETED
Standard SA95 U-Special, B-, V-Bessel, g-, r-Sloan; V-, R-Bessel, r-, i-, z-Sloan - airmass 1.2
Bellazzini_6K HVC196.09 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec COMPLETED
Bellazzini_6A HVC205.28 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec COMPLETED
Standard SA95 U-Special, B-, V-Bessel, g-, r-Sloan; V-, R-Bessel, r-, i-, z-Sloan - airmass 1.35
Bellazzini_6B HVC204.88 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec COMPLETED
Giallongo_44A MACSJ0717 U-Special, r-Sloan 30min, 30min within the specs
Bellazzini_6J Leo P g-Sloan, r-Sloan 10min+40sec, 10 min+40sec COMPLETED
Bellazzini_6Y HVC255.76 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec COMPLETED
Bellazzini_6U HVC250.16 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec COMPLETED
Bellazzini_6L HVC245.26 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec COMPLETED
Bellazzini_6C HVC277.25 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec COMPLETED
Bellazzini_6D HVC274.68 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec COMPLETED
Bellazzini_6M HVC298.95 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec COMPLETED
Bellazzini_6N HVC326.91 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec Executed, but out of seeing specs
Bellazzini_6O HVC28.09 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec Executed. Seeing to be cheked
Bellazzini_6P HVC353.41 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec COMPLETED
Bellazzini_6E HVC351.17 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec COMPLETED
Bellazzini_6Q HVC352.45 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec COMPLETED
Bellazzini_6F HVC356.81 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec COMPLETED
Bellazzini_6R HVC5.58 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec COMPLETED
Bellazzini_6G HVC13.59 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec COMPLETED

January 29-30

During the twilight we have thin cirrus passing, wind at 15 m/s and 40% humidity .
N.B. The seeing from the GCS is overestimated by 10% because the MODs guider is out of focus.
At the beginning of science integration the seeing is 1.4 on the GCS and 1.6 on the DIMM amd thin clouds are passing.
Thick clouds passing during the transit. The WFS was not sending corrections to the primary and observations are affected by astigmatism. The SA calls J. C. Guerra to try to increase the WFS exp time, but this is coupled with the guider. To do not miss the guiding we give it up.

Project_RUN Objects Filter ExpTime Notes
Standard G191b2b dual grating - airmass 1.1
Nascimbeni_DDT GJ3470b dual grating 4h 40 m Executed with some issues
Gilli Q1 J1048 Q1 red grating 1h 30m Executed out of seeing specs
Piconcelli C1 RBS 1273 red grating 5 m Executed out of seeing specs
Piconcelli C2 RBS 1124 red grating 5 m Executed out of seeing specs

January 30-31

At the beginning of the night we stay closed due to strong wind. Thick clouds are passing.
We stay close all the night long.
We run all MODS calibrations.

January 31- February 01

We stay closed due to strong wind (30 m/s) and temperature close to dew point.

February 01-02

We remain close because of temperature reaching the dew point.

February 02-03

At the beginning of the night we stay close because of dewpoint reaching ext temperature.
While we are waiting the TO starts the mirror ventilation.
At the beginning of scientific operations there are clouds passing, 60% humidity and no wind.
Since there are thick clouds we can't observe scientific targets: no collimation (the WFS can't sent corrections to collimate) + no guiding (the GCS can't see a R=12 star).
We try to collimate on a very bright star, this does not work and the TO closes the dome.
We stay closed the rest of the night.

Project_RUN Objects Filter ExpTime Notes
Standard Feige 34 red grating - 1.1

February 03-04

We stay close all the night long due to snow and dew-point temperature equal to the external one.

February 04-05

At twilight we thin cirrus pasassing from SE to NE, 2m/s blowing wind, no gusts and 20% humidity.
We have clouds passing, these become thick at ~8.30UT avoiding guiding and collimation.
The TO closes the dome at 9.30 due to clouds and increasing humidity.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Standard G191b2b dual grating - airmass 1.1
Vanzella_B macsj0647 dual grating 3h some clouds in the last 0.5h
Gilli_Q1 J1048 red grating 1h some clouds in the last 15min. MASK EXECTUTED

February 04-05

At sunset clouds, the dome stays closed. During the night we stay closed because of dew-point = ext temperature. After 3.00UT we could open, but with frost on the dome it is not safe.

-- EleanoraSani - 02 Feb 2014
Topic revision: r4 - 19 Feb 2014, EleanoraSani
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback